OpenStudy (anonymous):

4 more than a number is 16

OpenStudy (anonymous):

x+4=16 x=16-4=12

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

think about it this way... 4 MORE means 4 +... so number is not given which means it is an unknown value.. x.. so you have 4+x now... and it says it IS 16... which means =16. so you have 4 more (4+) than a number (x) is 16 (=16) put it al together and u get 4+x=16 so now you solve for x... what do u get?? or do u need to solve for x?

OpenStudy (anonymous):

A. 4 + 16 = n B. 4 × n = 16 C. n - 4 = 16 D. n + 4 = 16 2. Which of the followi

OpenStudy (anonymous):

D. n+4=16 cuz the unknown variable can be anything. do u understand?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yup :) just remember to use key words.

OpenStudy (anonymous):

7 less than a number is 16 A. B. C. D.

OpenStudy (anonymous):

do the same thing... what do u think?

OpenStudy (anonymous):

**same thing as the last one

OpenStudy (anonymous):

x-7=16